OpenStudy (anonymous):

Square root 3 * cot(theta) - 1 =0

OpenStudy (anonymous):

\[\sqrt3*\cot \theta -1 = 0\] i hope this is what you mean

OpenStudy (anonymous):

Yea I need to find all solutions

OpenStudy (anonymous):

Unsure on how to do so

OpenStudy (anonymous):

then sqrt 3 cot theta = 1 and cot theta = 1/ sqrt 3 then theta would equals cot inverse of 1/ sqrt 3

OpenStudy (anonymous):

got it till now?

OpenStudy (anonymous):

How would I answer in terms of radians. I remember something about a K

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4ff530c1e4b01c7be8c84e86 see it once

OpenStudy (anonymous):

So am I right. \[\Theta = k \pi + \frac{ \pi }{ 4\sqrt{3} }\]

OpenStudy (anonymous):

Hmm :)

OpenStudy (anonymous):

wait